# OlympiadBench: A Challenging Benchmark for Promoting AGI with Olympiad-Level Bilingual Multimodal Scientific Problems

### v1

@@ Appendix A Dataset Details
Table 5: Summary of Problems in Maths and Physics Competitions, the full name of each acronym is given in Table 6
@@ A.1 Data Sources
    Gaokao Mock Questions for Mathematics and Physics  [^11] Due to anonymity of this submission, we can not add concrete source in this version.. Given that the resolution of Olympiad-level problems typically necessitates models with substantial parameter sizes, we also incorporate Gaokao simulation problems to evaluate smaller models’ capabilities in answering free-form mathematics and physics questions.
@@ A.3 Example of Progressive Problem in Physics
Figures 7 and 8 present a sequential challenge from the International Physics Olympiad (IPhO) 2021, illustrating the intricacies of progressive problem-solving in a competitive context. This particular problem set exemplifies a common trait in advanced physics competitions: the dependency of many questions on the solutions and materials of preceding ones. These dependencies are sometimes explicit, but most are implicit, weaving a complex web of interconnected knowledge and reasoning.
@@ Appendix C Additional Analysis and Examples
C.1 Detailed Description of the Error Types in GPT-4V’s solving process
@@
    Question Misunderstanding
@@
    Value Calculation Error
@@
    Expression Calculation Error
@@
    Logical Reasoning / Induction Error / Conceptual Confusion
@@
    Introducing Unnecessary variables or concepts
@@
    Conclusion Hallucination
@@
    Unfinished Answering: sometimes GPT-4V says the question have confliction in settings, or degenerates after some tokens.
@@
    Insufficient Classification Discussions
@@
    Incorrect Judging: Calls for future work of automatically deciding required precision of the answer, or automatically judging expressions such as $a\sqrt{b}$ and $\sqrt{a^{2}b}$ with $a\geq 0$.
C.2 Other Observation for Open-Ended Problems
@@
Sequence 273
Trigonometric Functions 244
Set Theory 25
@@
Elementary Functions 167
Inequality 139
PC&PE 95
Probability and Statistics 865
Plane Geometry 831
Solid Geometry 1375
@@
COMP Math Combinatorics 406
Algebra 567
Number Theory 295
Geometry 544
CEE&COMP Physics Mechanics 1040
Electromagnetism 756
Thermodynamics 257
Optics 157
Modern Physics 220

### v2

@@ 7 Conclusion
Acknowledgements
This work is supported by the National Key R&D Program of China (No.2022ZD0116312), National Natural Science Foundation of China (No. 62236004), Quan Cheng Laboratory (Grant No. QCLZD202301) and Institute Guo Qiang at Tsinghua University.
@@ Appendix A Dataset Details
Table 5: Summary of Problems in Math and Physics Competitions, with full acronyms listed in the Table 6
@@ A.1 Data Sources
    Gaokao Mock Questions for Mathematics and Physics. Given that the resolution of Olympiad-level problems typically necessitates models with substantial parameter sizes, we also incorporate Gaokao simulation problems to evaluate smaller models’ capabilities in answering free-form mathematics and physics questions.
@@ A.3 Example of Progressive Problem in Physics
Figures 10 and 11 present a sequential challenge from the International Physics Olympiad (IPhO) 2021, illustrating the intricacies of progressive problem-solving in a competitive context. This particular problem set exemplifies a common trait in advanced physics competitions: the dependency of many questions on the solutions and materials of preceding ones. These dependencies are sometimes explicit, but most are implicit, weaving a complex web of interconnected knowledge and reasoning.
@@ Appendix C Additional Analysis and Examples
C.1 Performance analysis of GPT-4V
We analyzed GPT-4V’s performance (accuracy on open-ended problems) on different knowledge points based on the knowledge point labels in OlympiadBench, the results can be found at Figure 5.
For Math problems, GPT-4V has poor performance in geometry, with the lowest scoring knowledge points being almost exclusively geometry-related. This may show the need of improving the ability of understanding and imaging plane or 3d situations. GPT-4V also performs poorly on knowledge points that are more computationally intensive such as conic curves; and struggles to give a complete and comprehensive classification discussion, therefore prone to making mistakes on combinatorial problems. However, GPT-4V is stronger in knowledge related to derivatives and complex numbers.
As for Physics problems, none of the knowledge points surpass an accuracy of 16%, and GPT-4V struggles more in thermodynamics and mechanics.
Figure 5: An analysis of GPT-4V’s accuracy on different knowledge points, blue for Math and red for Physics.
C.2 Detailed Description of the Error Types in GPT-4V’s solving or proving process
@@
    Question Misunderstanding: GPT-4V sometimes misunderstands the intention or settings of the question.
@@
    Value Calculation Error: GPT-4V make simple calculation mistakes sometimes, such as outputting $\frac{b}{2}+7=\frac{b+7}{2}$, these mistakes appears more in Chinese and Math contents.
@@
    Expression Calculation Error: Similar to value calculation error, but happens when transforming between two expressions.
@@
    Logical Reasoning / Induction Error / Conceptual Confusion: GPT-4V sometimes makes false reasoning or induction, as well as encounters conceptual confusion (see Figure 7 for example).
@@
    Introducing Unnecessary variables or concepts: GPT-4V sometimes suddenly introduce variables or try to use concepts that have no contribution to solving the problem, which not only makes the output longer, but also may confuse GPT-4V itself and leads to incorrect output.
@@
    Conclusion Hallucination: GPT-4V sometimes hallucinates for a conclusion that is not reached in former output, or hallucinates a theorem that does not really exist (for example, when solving geometric proving problem, GPT-4V always mention "The Power Theorem", which does not exist, and all the proof thereafter will lost their logic).
@@
    Unfinished Answering: GPT-4V sometimes says the question have confliction in settings (which is not true), or degenerates after some tokens.
@@
    Insufficient Classification Discussions: When doing classification discussion, GPT-4V may miss some possible situation, or have overlapped discussion (see Figure 6 for example).
@@
    Incorrect Judging: Sometimes GPT-4V gives the right answer, but is judged as incorrect due to the limitation of the automated scoring system: One important problem is that many problems, especially Physics problems, accept answers that fall in a specific range due to rounding up, rather than a fixed numerical answer, so a precision is needed for automatically calculating the range, which was not given in many cases. Using relative precision (such as a percentage of 1%), may leads to accepting answers with unacceptable error when the origin answer is big, so we manually decide the precision for OlympiadBench problems if there is no official precision given, which may leads to incorrect judging. For example, in a Physics question (Physics-En_COMP#995), the precision is manually set as $11e$ 1, with the official answer being $\sqrt{2}\times 3970\approx 5614.43$. GPT-4V calculates the answer as $\sqrt{2}\times 3970\approx 1.414\times 3970\approx 5615.88$, which falls out of the acceptable range, but is actually correct. Another problem is the current automated scoring system can’t judge the equality between expressions such as $a\sqrt{b}$ and $\sqrt{a^{2}b}$ with $a\geq 0$. These flaws call for future work such as automatically deciding required precision of the answer, and automatically judging equality between expressions.
Figure 6: An example of GPT-4V making insufficient classification discussions in a combinatorial problem. The red text is the wrong part of GPT-4V’s answer, and the purple text is our further illustration.
Figure 7: An example of GPT-4V getting confused between the concept of "vertical to the plane" and "vertical(to the ground)".
C.3 Other Observation for Open-Ended Problems
@@
    GPT-4V may success in giving correct overall idea, but fail in calculation (such as solving quadratic equations with extra negative signs), which leads to a wrong answer.
  4. 4.
@@
Sequence 258
Trigonometric Functions 236
Set Theory 22
@@
Elementary Functions 158
Inequality 138
PC&PE 80
Probability and Statistics 758
Plane Geometry 790
Solid Geometry 1322
@@
COMP Math Combinatorics 369
Algebra 525
Number Theory 256
Geometry 535
CEE&COMP Physics Mechanics 1010
Electromagnetism 714
Thermodynamics 248
Optics 153
Modern Physics 209